• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ramesh

    Log entry time 00:16:07 on April27,2008

    Entry number 234929

    keyword=Shift summary

    Kinematics K.

    Beam energy: 119.56 MeV
    Asking current: 3.5 uA


    Left HRS:
    angle: 14.503 deg
    momentum: 0.997 GeV/c

    Right HRS:
    angle: 104.998 deg
    momentum: 0.4596 GeV/c




    16:00 Shift starts with run 4017

    16:25 Could not take a single datum in the run 4019. Found BB HV tripping. After repairing the BB HV, had to reboot ROC5. Had to reboot ROC9 and ROC10 to be able to make the DAQ running.

    16:48 Good to run the DAQ. Start of run 4020-

    19:06 BB HV trip.

    21:08 Changed target to Al dummy (not the Al dumy holy one) for
    the dummy target run for an hour, without changing any running
    conditions compared to production run.

    22:43 Started regular production runs on LH2 6cm.


    Run summary:

    Run No Events Target Comments

    4017 1M LH2
    4018 1M LH2
    4019 Junk
    4020 1M LH2 End of Run failed, good run
    4021-404025 1M LH2
    4026 LH2 junk
    4027-4030 1M LH2
    4031 0.5M LH2 probably good, BB HV tripped
    4032-4037 1M LH2

    4038-4040 1M Al dummy
    4041 0.5M Al dummy
    4042-4049 1M LH2